# Codes Correcting Two Bursts of Exactly $b$ Deletions

### v4

@@ III Bounds
Since $m=n/b-1$, by definition of $\mathcal{C}^{\prime}$, we have
$\left|\mathcal{C}^{\prime}\right|$ =|$=\left|\left\{\mathbf{v}\in\Sigma_{q}^{m+1}:r(\mathbf{v})\leq t+1\right\}\right|^{b}$ (10)
@@
$=\left(\sum_{j=1}^{m+1}\left|\left\{\mathbf{v}\in\Sigma_{q}^{m+1}:r(\mathbf{v})=j\right\}\right|\right)^{b}$
=$\overset{(a)}{=}\left(q\sum_{j=1}^{t+1}\binom{m}{j-1}(q-1)^{j-1}\right)^{b}=\left(q\sum_{j=0}^{t}\binom{m}{j}(q-1)^{j}\right)^{b}$
where (a) follows from the well-known result (see the proof of [41, Theorem 3.1]):
@@
Since $t=(1-1/q-\epsilon)m$, we have $t/m=1-1/q-\epsilon\leq 1-1/q$. Then it follows from [42, Proposition 3.3.3] that
@@
If $\frac{\log n}{n}<\frac{\log q}{12}f(q)^{2}$, then $\epsilon<f(q)$. Now by Lemma .1, we have
@@
$\left|\mathcal{C}_{2}\right|\leq\left|\mathcal{C}^{\prime}\right|\leq q^{b+(n-b)(1-\epsilon^{2}/4)}\overset{(a)}{=}q^{b+(n-b)\left(1-\frac{3\log n}{n\log q}\right)}=\frac{q^{n}}{n^{3}}\left(\sqrt[n]{n}\right)^{3b}\overset{(b)}{<}(1.121)^{3b}\frac{q^{n}}{n^{3}},$ (12)
where (a) follows from selecting $\epsilon=\sqrt{\frac{12\log n}{n\log q}}$ and (b) follows from the fact that $n]{n}<1.121$ when $n\geq 30$. We conclude the proof by combining Equations 9 and 12. ∎
@@
$2\log n+(2b-2)\log q+2\log(q-1)-2\log b-2\leq\rho(\mathcal{C})\leq 4\log n+2b\log q,$
where $\mathcal{C}\subseteq\Sigma_{q}^{n}$ is a $2$-$b$-burst-deletion correcting code with maximum size.
@@ IV Codes For Correcting Two $b$-Burst-Deletions
In this section, we construct $q$-ary $2$-$b$-burst-deletion correcting codes with redundancy at most $5\log n+O(\log\log n)$, for any $q\geq 2$ and $b>1$. Our idea is to first locate positions of deletions in short intervals (which is accomplished in Section IV-A, Lemma IV.5) and then correct errors in these intervals (which is accomplished in Sections IV-B and IV-C).
@@ IV-A Approximately determine positions of deletions
[19, Theorem 7] Suppose $d\geq 7$. There exists a function $\eta:\Sigma_{2}^{n}\rightarrow\Sigma_{2}^{4\log n+10\log\log n+O_{d}(1)}$, computable in linear time, such that for any $d$-regular sequence $\mathbf{x}\in\Sigma_{2}^{n}$, given $\eta\left(\mathbf{x}\right)$ and any $\mathbf{y}\in\mathcal{B}_{2}\left(\mathbf{x}\right)$, one can uniquely and efficiently recover $\mathbf{x}$. Here, $O_{d}(1)$ denotes a constant depending only on $d$.
If the first row $A(\mathbf{x})_{1}$ is regular, then by correcting the two deletions in $A(\mathbf{x})_{1}$, we can approximate the location of the two deletions, as ensured by the following lemma.
@@
[21, Lemma 9] Suppose that $\mathbf{x}\in\Sigma_{2}^{n}$ is $d$-regular and $\mathbf{x}^{\prime}\in\Sigma_{2}^{n-2}$ is obtained from $\mathbf{x}$ by deleting two symbols $x_{i_{1}}$ and $x_{i_{2}}$. Then we can
@@ 2. $(2)$
We briefly explain why Lemma IV.3 is helpful. Suppose $\mathbf{x}\in\Sigma_{2}^{n}$ and $\mathbf{x}^{\prime}\in\mathcal{B}_{2}^{b}(\mathbf{x})$. Let $A(\mathbf{x})=A(\mathbf{x},n,b)$ and $A(\mathbf{x}^{\prime})=A(\mathbf{x}^{\prime},n-2b,b)$. Then $A(\mathbf{x}^{\prime})_{i}\in\mathcal{B}_{2}(A(\mathbf{x})_{i})$ for all $1\leq i\leq 2$ If $A(\mathbf{x})_{1}$ is $d$-regular (where $d\geq 7$) and $\eta(A(\mathbf{x})_{1})$ is given, Lemma IV.2 ensures that we can decode $A(\mathbf{x})_{1}$ from $A(\mathbf{x}^{\prime})_{1}$ and Lemma IV.3 ensures that error positions in $A(\mathbf{x})_{1}$ can be approximately determined in one or two short intervals. Then by II.2, we can approximate error positions in $A(\mathbf{x})_{i}$ within one or two short intervals, for all $2\leq i\leq b$.
However, the above analysis is valid only when $\mathbf{x}$ is a binary sequence. Since we deal with binary and non-binary cases in a unified way, it is convenient to generalize the notion of regularity to general alphabets. To that end, we associate with any $q$-ary sequence a binary sequence.
Let $q\geq 2$. For any $x\in\Sigma_{q}$, we can uniquely write it as $x=\left\lceil q/2\right\rceil u_{x}+v_{x}$, where $u_{x}\in\{0,1\}$ and $0\leq v_{x}<\left\lceil q/2\right\rceil$. Then for a sequence $\mathbf{x}\in\Sigma_{q}^{n}$, define $u(\mathbf{x})\triangleq u_{x_{1}}\cdots u_{x_{n}}$, i.e., $u(\mathbf{x})_{i}=u_{x_{i}}$ for all $1\leq i\leq n$. Clearly, $u(\mathbf{x})\in\Sigma_{2}^{n}$ and when $q=2$, we have $u(\mathbf{x})=\mathbf{x}$. It should be noted that this decomposition of $x\in\Sigma_{q}$ into $u_{x}$ and $v_{x}$ was also applied in [23, Section V] to construct a single-burst-deletion of variable length.
@@
[…]
@@
  • If $q$ is odd, then $\left|\mathcal{R}_{q,n,d}\right|\geq q^{n-1}$ for all $d\geq 10$ and $\left\lfloor\frac{d}{2}\log n\right\rfloor n^{-1-\frac{d}{2}\log(0.87)}\geq\frac{200q}{87(q-1)}$. Note that when $d\geq 10$, we have $-1-\frac{d}{2}\log(0.87)>0$ and hence this condition could be satisfied when $n$ is sufficiently large. In particular, when $d=10$, it suffices to require $\left\lfloor 5\log n\right\rfloor n^{-1-5\log(0.87)}\geq\frac{200q}{87(q-1)}$.
@@
Suppose first that $q$ is even. Any $x\in\Sigma_{q}$ can be uniquely represented as $x=\frac{q}{2}u_{x}+v_{x}$, where $u_{x}\in\{0,1\}$ and $20\leq v_{x}<q/$ 2. On the other hand, when $q$ is even, we have $\frac{q}{2}u_{x}+v_{x}\in\Sigma_{q}$ for any $u_{x}\in\{0,1\}$ and $20\leq v_{x}<q/$ 2. Then it follows from Lemma IV.1 that $\left|\mathcal{R}_{q,n,d}\right|\geq\left|\mathcal{R}_{2,n,d}\right|(q/2)^{n}=q^{n}/2\geq q^{n-1}$.
@@
$S_{m}^{L}=\left\{\mathbf{v}\in\Sigma_{q}^{m}:\nexists 1\leq i<m\text{ such that }0\leq v_{i},v_{i+1}\leq(q-1)/2\right\},$
$S_{m}^{U}=\left\{\mathbf{v}\in\Sigma_{q}^{m}:\nexists 1\leq i<m\text{ such that }(q+1)/2\leq v_{i},v_{i+1}<q\right\}.$
We next prove by induction on $m$ that $\left|S_{m}^{L}\right|<(0.87q)^{m}$ for all $m\geq 2$. By the inclusion-exclusion principle, it is easy to verify that $\left|S_{2}^{L}\right|=q^{2}-\left(\frac{q+1}{2}\right)^{2}=\frac{3}{4}q^{2}-\frac{1}{2}q-\frac{1}{4}<0.75q^{2}<(0.87q)^{2}$ and $\left|S_{3}^{L}\right|=q^{3}-2q\left(\frac{q+1}{2}\right)^{2}+\left(\frac{q+1}{2}\right)^{3}=\frac{5}{8}q^{3}-\frac{5}{8}q^{2}-\frac{1}{8}q+\frac{1}{8}<0.625q^{3}<(0.87q)^{3}$. Now suppose that $m\geq 4$ and the conclusion is proved for all $m^{\prime}\leq m-2$. For $\mathbf{v}\in S_{m}^{L}$, it must holds that $v_{1}v_{2}\in S_{2}^{L}$ and $v_{3}\cdots v_{m}\in S_{m-2}^{L}$. So it follows that $\left|S_{m}^{L}\right|\leq\left|S_{2}^{L}\right|\left|S_{m-2}^{L}\right|<(0.87q)^{2}(0.87q)^{m-2}=(0.87q)^{m}$.
@@
$Q=\left\{\mathbf{v}\in\Sigma_{q}^{m}:\exists i,j\text{ such that }v_{i},v_{i+1}<(q+1)/2\text{ and }v_{j}.v_{j+1}\geq(q+1)/2\right\}.$
Then we have $\left|Q\right|\geq q^{m}-\left|S_{m}^{L}\right|-\left|S_{m}^{U}\right|>q^{m}-2(0.87q)^{m}$. Now let $k=\left\lfloor n/m\right\rfloor$ and define
=$\mathcal{R}^{\prime}=\left\{\mathbf{x}=\mathbf{x}^{(1)}\cdots\mathbf{x}^{(k)}\mathbf{x}^{(k+1)}\in\Sigma_{q}^{n}:\begin{array}[]{c}\mathbf{x}^{(i)}\in Q,\forall 1\leq i\leq k\\ \mathbf{x}^{(k+1)}\in\Sigma_{q}^{n-k\left\lfloor n/m\right\rfloor}.\end{array}\right\}$
It is straightforward to see that $\mathcal{R}^{\prime}\subseteq\mathcal{R}_{q,n,d}$ and thus
@@
Notations: In the rest of this section, for any $q$-ary sequence $\mathbf{x}$, denote $A(\mathbf{x})=A(\mathbf{x},\left|\mathbf{x}\right|,b)$ and $U(\mathbf{x})=A(u(\mathbf{x}),\left|\mathbf{x}\right|,b)$. Denote the $i$-th row of $A(\mathbf{x})$ and $U(\mathbf{x})$ by $A(\mathbf{x})_{i}$ and $U(\mathbf{x})_{i}$, respectively. It is clear that $U(\mathbf{x})_{i}=u\left(A(\mathbf{x})_{i}\right)$ for all $1\leq i\leq b$. Let
$I_{n,b}\triangleq\left\{1+(j-1)b:1\leq j\leq n/b\right\},$
which is a subset of $[n]$. Then we have $A(\mathbf{x})_{1}=\mathbf{x}_{I_{n,b}}$ and $U(\mathbf{x})_{1}=u(\mathbf{x})_{I_{n,b}}$.
@@
Recall that $r\left(U(\mathbf{x})_{1}\right)$ is the number of runs in $U(\mathbf{x})_{1}$. For $1\leq j\leq r\left(U(\mathbf{x})_{1}\right)$, let $U(\mathbf{x})_{1,I_{j}}$ be the $j$-th run. Furthermore, suppose $I_{j}=\left[p_{j-1}+1,p_{j}\right]$, where $p_{0}=0$, $p_{r\left(U(\mathbf{x})_{1}\right)}=n/b$ and $p_{j-1}<p_{j}$ for all $j$. By definition, $I_{j}$ is an interval in $[n/b]$. For all $j\geq 1$, we associate with $I_{j}$ an interval in $[n]$: $I_{j}^{\prime}\triangleq\left[p_{j-1}b+1,p_{j}b\right]$. In addition, define $I_{j}^{L}\triangleq\left[(p_{j-1}-1)b+2,p_{j-1}b\right]\cap[n]$. In other words, $I_{j}^{L}$ is the interval of length at most $b-1$ in $[n]$ to the left of $I_{j}^{\prime}$. See Figure 1 for an illustration. Note that $I_{j}$, $I_{j}^{\prime}$ and $I_{j}^{L}$ are dependent on $\mathbf{x}$. We omit $\mathbf{x}$ for simpler notations.
@@
Let $q\geq 2$, $b>1$ and $n>2b$. Suppose $\mathbf{x}\in\Sigma_{q}^{n}$ and $\mathbf{x}^{\prime}=\mathbf{x}_{[n]\setminus(D_{1}\cup D_{2})}$, where $D_{1},D_{2}\subseteq[n]$ are two disjoint intervals of length $b$. Suppose that $A(\mathbf{x})_{1}$ is $d$-regular, that is, $U(\mathbf{x})_{1}$ is $d$-regular. Let $\eta(\cdot)$ be the function defined in Lemma IV.2. Then given $\eta\left(U(\mathbf{x})_{1}\right)$, we can

### v5

@@ III Bounds
Since m=n/b1m=n/b-1, by definition of $\mathcal{C}^{\prime}$, we have
$\left|\mathcal{C}^{\prime}\right|$ $=\left|\left\{\mathbf{v}\in\Sigma_{q}^{m+1}:r(\mathbf{v})\leq t+1\right\}\right|^{b}$ (10)
@@
$=\left(\sum_{j=1}^{t+1}\left|\left\{\mathbf{v}\in\Sigma_{q}^{m+1}:r(\mathbf{v})=j\right\}\right|\right)^{b}$
$\overset{(a)}{=}\left(q\sum_{j=1}^{t+1}\binom{m}{j-1}(q-1)^{j-1}\right)^{b}$
$=\left(q\sum_{j=0}^{t}\binom{m}{j}(q-1)^{j}\right)^{b}$
where (a) follows from the well-known result (see the proof of [42, Theorem 3.1]):
@@
Since $t=(1-1/q-\epsilon)m$, we have $t/m=1-1/q-\epsilon\leq 1-1/q$. Then it follows from [43, Proposition 3.3.3] that
@@
If $\frac{\log n}{n}<\frac{\log q}{12}f(q)^{2}$, then $\epsilon<f(q)$. Now by Lemma A.1, we have
@@
$\left|\mathcal{C}_{2}\right|\leq\left|\mathcal{C}^{\prime}\right|\leq q^{b+(n-b)(1-\epsilon^{2}/4)}$ $\overset{(a)}{=}q^{b+(n-b)\left(1-\frac{3\log n}{n\log q}\right)}$ (12)
$=\frac{q^{n}}{n^{3}}\left(\sqrt[n]{n}\right)^{3b}$
[…]
where (a) follows from selecting $\epsilon=\sqrt{\frac{12\log n}{n\log q}}$ and (b) follows from the fact that $\sqrt[n]{n}<1.121$ when $n\geq 30$. We conclude the proof by combining Equations 9 and 12. ∎
@@
$2\log n+\log\left(\frac{q^{2b-2}(q-1)^{2}}{4b^{2}}\right)\leq\rho(\mathcal{C})\leq 4\log n+2b\log q,$
where $\mathcal{C}\subseteq\Sigma_{q}^{n}$ is a 22-$b$-burst-deletion correcting code with maximum size.
Proof:
Firstly, the upper bound follows straightforward from Theorem III.1 just by noticing that $\binom{n}{2}^{2}\leq n^{4}$. Since $n\geq\frac{(1.21)^{3b}q^{2b-2}(q-1)^{2}}{b^{2}}$, we have $\left(\frac{3b^{2}}{q^{2b-2}(q-1)^{2}}+\frac{(1.121)^{3b}}{n}\right)\frac{q^{n}}{n^{2}}\leq\frac{4b^{2}}{q^{2b-2}(q-1)^{2}}\frac{q^{n}}{n^{2}}$. Now the lower bound follows from Theorem III.2. ∎
@@ IV Codes For Correcting Two $b$-Burst-Deletions
In this section, we construct $q$-ary 22-$b$-burst-deletion correcting codes with redundancy at most $5\log n+O(\log\log n)$, for any $q\geq 2$ and $b>1$. Our idea is to first locate positions of deletions in short intervals (which is accomplished in Section IV-A, Lemma IV.5) and then correct errors in these intervals (which is accomplished in Sections IV-B and IV-C).
@@ IV-A Approximately determine positions of deletions
The next lemma ensures that we can efficiently correct two deletions in $d$-regular sequences.
@@
[19, Theorem 7] Suppose $d\geq 7$. There exists a function $\eta:\Sigma_{2}^{n}\rightarrow\Sigma_{2}^{4\log n+10\log\log n+O_{d}(1)}$, computable in linear time, such that for any $d$-regular sequence $\mathbf{x}\in\Sigma_{2}^{n}$, given $\eta\left(\mathbf{x}\right)$ and any $\mathbf{x}^{\prime}\in\mathcal{B}_{2}\left(\mathbf{x}\right)$, one can uniquely and efficiently recover $\mathbf{x}$. Here, $O_{d}(1)$ denotes a constant depending only on $d$.
Furthermore, once $\mathbf{x}$ is recovered, locations of the two deletions in $\mathbf{x}$ can be approximated, as ensured by the following lemma.
@@
[21, Lemma 9] Suppose that $\mathbf{x}\in\Sigma_{2}^{n}$ is $d$-regular and $\mathbf{x}^{\prime}\in\Sigma_{2}^{n-2}$ is obtained from $\mathbf{x}$ by deleting two symbols $x_{i_{1}}$ and $x_{i_{2}}$. When given $\mathbf{x}$ and $\mathbf{x}^{\prime}$, we can
@@ 2. $(2)$
A proof of Lemma IV.3 is given in [21]. Here we show the intuition behind this lemma.
Example IV.1
  1. $(1)$
    Let $\mathbf{x}=000111$ and $\mathbf{x}^{\prime}=0011$. Then $\mathbf{x}^{\prime}$ is obtained from $\mathbf{x}$ by deleting one bit in the run $000$ and one bit in the run $111$. This corresponds to case (1) in Lemma IV.3.
  2. $(2)$
    $\mathbf{x}=10010101110$ and $\mathbf{x}^{\prime}=100101110$. Then $\mathbf{x}^{\prime}$ can be obtained from $\mathbf{x}$ by deleting $x_{2}$ and $x_{4}$, or by deleting $x_{3}$ and $x_{4}$, or by deleting $x_{4}$ and $x_{5}$, or by deleting $x_{7}$ and $x_{8}$, or by deleting $x_{7}$ and $x_{9}$. Therefore, we can only locate error positions in the substring $001010111$, which is the concatenation of a run $00$, an alternating substring $1010$ and a run $111$. If $\mathbf{x}\in\Sigma_{2}^{n}$ is $d$-regular, then each run has length at most $d\log n$ and each alternating substring has length at most $d\log n$. This is why in case (2) of Lemma IV.3 we can locate error positions in an interval of length at most $3d\log n$.
Now we briefly explain why Lemma IV.3 is helpful. Suppose $\mathbf{x}\in\Sigma_{2}^{n}$ and $\mathbf{x}^{\prime}\in\mathcal{B}_{2}^{b}(\mathbf{x})$. Let $A(\mathbf{x})=A(\mathbf{x},b)$ and $A(\mathbf{x}^{\prime})=A(\mathbf{x}^{\prime},b)$ be the matrix representations of $\mathbf{x}$ and $\mathbf{x}^{\prime}$, respectively. If $A(\mathbf{x})_{1}$ is $d$-regular (where $d\geq 7$) and $\eta(A(\mathbf{x})_{1})$ is given, Lemma IV.2 ensures that we can decode $A(\mathbf{x})_{1}$ from $A(\mathbf{x}^{\prime})_{1}$ and Lemma IV.3 ensures that error positions in $A(\mathbf{x})_{1}$ can be approximately determined in one or two short intervals. By II.2, this further reveals information of error positions in remaining rows of $A(\mathbf{x})$, or equivalently, error positions in $\mathbf{x}$. Detailed analysis will be given Lemma IV.5. Before that, it is convenient to generalize the notion of regularity to general alphabets. To that end, we associate with any $q$-ary sequence a binary sequence.
Let $q\geq 2$. For any $x\in\Sigma_{q}$, we can uniquely write it as $x=\left\lceil q/2\right\rceil u_{x}+v_{x}$, where $u_{x}\in\{0,1\}$ and $0\leq v_{x}<\left\lceil q/2\right\rceil$. Then for a sequence $\mathbf{x}\in\Sigma_{q}^{n}$, define $u(\mathbf{x})\triangleq u_{x_{1}}\cdots u_{x_{n}}$, i.e., $u(\mathbf{x})_{i}=u_{x_{i}}$ for all $1\leq i\leq n$. We call $u(\mathbf{x})$ the binary sequence associated with $\mathbf{x}$. Clearly, $u(\mathbf{x})\in\Sigma_{2}^{n}$ and when $q=2$, we have $u(\mathbf{x})=\mathbf{x}$. It should be noted that this decomposition of $x\in\Sigma_{q}$ into $u_{x}$ and $v_{x}$ was also applied in [23, Section V] to construct a single-burst-deletion of variable length.
Example IV.2
Let $q=3$. We have that $\left\lceil q/2\right\rceil=2$. If $x=1$, then $u_{x}=0$ and $v_{x}=1$. If $x=2$, then $u_{1}=1$ and $v_{x}=0$.
@@
Let $\mathcal{R}_{q,n,d}\triangleq\left\{\mathbf{x}\in\Sigma_{q}^{n}:\mathbf{x}\text{ is }d\text{-regular}\right\}$. To estimate a lower bound on our code size in Theorem IV.1, we need a lower bound on $\left|\mathcal{R}_{q,n,d}\right|$.
@@
  • If $q$ is odd, then $\left|\mathcal{R}_{q,n,d}\right|\geq q^{n-1}$ for all $d\geq 10$ and $\left\lfloor\frac{d}{2}\log n\right\rfloor n^{-1-\frac{d}{2}\log(0.87)}\geq\frac{200q}{87(q-1)}$. Note that when $d\geq 10$, we have $1-\frac{d}{2}\log(0.87)>0$ and hence this condition could be satisfied when $n$ is sufficiently large. In particular, when $d=10$, it suffices to require $\left\lfloor 5\log n\right\rfloor n^{-1-5\log(0.87)}\geq\frac{200q}{87(q-1)}$.
@@
Suppose first that $q$ is even. Any $x\in\Sigma_{q}$ can be uniquely represented as $x=\frac{q}{2}u_{x}+v_{x}$, where $u_{x}\in\{0,1\}$ and $0\leq v_{x}<q/2$. On the other hand, when $q$ is even, we have $\frac{q}{2}u_{x}+v_{x}\in\Sigma_{q}$ for any $u_{x}\in\{0,1\}$ and $0\leq v_{x}<q/2$. Then it follows from Lemma IV.1 that $\left|\mathcal{R}_{q,n,d}\right|\geq\left|\mathcal{R}_{2,n,d}\right|(q/2)^{n}=q^{n}/2\geq q^{n-1}$.
@@
$S_{m}^{L}=\left\{\mathbf{v}\in\Sigma_{q}^{m}:\begin{array}[]{c}\nexists 1\leq i<m\text{ such that}\\ 0\leq v_{i},v_{i+1}\leq(q-1)/2\end{array}\right\},$
$S_{m}^{U}=\left\{\mathbf{v}\in\Sigma_{q}^{m}:\begin{array}[]{c}\nexists 1\leq i<m\text{ such that}\\ (q+1)/2\leq v_{i},v_{i+1}<q\end{array}\right\}.$
We next prove by induction on $m$ that $\left|S_{m}^{L}\right|<(0.87q)^{m}$ for all $m\geq 2$. By the inclusion-exclusion principle, it is easy to verify that $\left|S_{2}^{L}\right|=q^{2}-\left(\frac{q+1}{2}\right)^{2}=\frac{3}{4}q^{2}-\frac{1}{2}q-\frac{1}{4}<0.75q^{2}<(0.87q)^{2}$ and $\left|S_{3}^{L}\right|=q^{3}-2q\left(\frac{q+1}{2}\right)^{2}+\left(\frac{q+1}{2}\right)^{3}=\frac{5}{8}q^{3}-\frac{5}{8}q^{2}-\frac{1}{8}q+\frac{1}{8}<0.625q^{3}<(0.87q)^{3}$. Now suppose that $m\geq 4$ and the conclusion is proved for all $m^{\prime}\leq m-2$. For $\mathbf{v}\in S_{m}^{L}$, it must hold that $v_{1}v_{2}\in S_{2}^{L}$ and $v_{3}\cdots v_{m}\in S_{m-2}^{L}$. So it follows that $\left|S_{m}^{L}\right|\leq\left|S_{2}^{L}\right|\left|S_{m-2}^{L}\right|<(0.87q)^{2}(0.87q)^{m-2}=(0.87q)^{m}$.
@@
$Q=\left\{\mathbf{v}\in\Sigma_{q}^{m}:\begin{array}[]{c}\exists i,j\text{ such that }v_{i},v_{i+1}<(q+1)/2\\ \text{and }v_{j}.v_{j+1}\geq(q+1)/2\end{array}\right\}.$
In other words, $Q$ is the set of length-$m$ sequences which contain two consecutive coordinates smaller than $\left\lceil q/2\right\rceil$ and two consecutive coordinates no less than $\left\lceil q/2\right\rceil$. Then we have $\left|Q\right|\geq q^{m}-\left|S_{m}^{L}\right|-\left|S_{m}^{U}\right|>q^{m}-2(0.87q)^{m}$. Now let $k=\left\lfloor n/m\right\rfloor$ and define $\mathcal{R}^{\prime}$ to be the following set
$\left\{\mathbf{x}=\mathbf{x}^{(1)}\cdots\mathbf{x}^{(k)}\mathbf{x}^{(k+1)}\in\Sigma_{q}^{n}:\begin{array}[]{c}\mathbf{x}^{(i)}\in Q,\forall 1\leq i\leq k\\ \mathbf{x}^{(k+1)}\in\Sigma_{q}^{n-km}\end{array}\right\}.$
It is easy to see that for any $\mathbf{x}\in\mathcal{R}^{\prime}$, any substring of $\mathbf{x}$ of length at least $d\log n$ must contain some $\mathbf{x}^{(i)}$, where $1\leq i\leq k$. Therefore, $\mathbf{x}$ is $d$-regular and hence $\mathcal{R}^{\prime}\subseteq\mathcal{R}_{q,n,d}$. This implies that
@@
In Appendix B, we will discuss how to encode a sequence into a $d$-regular sequence.
Now we return to the aim of this subsection: approximately determine error positions. Let $\mathbf{x}\in\Sigma_{q}^{n}$ and $\mathbf{x}^{\prime}$ be obtained from $\mathbf{x}$ by a 22-$b$-burst-deletion. Let $u(\mathbf{x})$ and $u(\mathbf{x}^{\prime})$ be the binary sequences associated with $\mathbf{x}$ and $\mathbf{x}^{\prime}$, respectively (see the paragraph prior to Example IV.2). Then we have $u(\mathbf{x}^{\prime})\in\mathcal{B}_{2}^{b}\left(u(\mathbf{x})\right)$. By the relationship between $\mathbf{x}$ and $u(\mathbf{x})$, it suffices to locate the 22-$b$-burst-deletion in $u(\mathbf{x})$. In the rest of this section, let $U(\mathbf{x})=A(u(\mathbf{x}),b)$ be the matrix representation of $u(\mathbf{x})$. Denote the $i$-th row of $U(\mathbf{x})$ by $U(\mathbf{x})_{i}$. Similarly, we can define $U(\mathbf{x}^{\prime})$ and $U(\mathbf{x}^{\prime})_{i}$.
Since $u(\mathbf{x}^{\prime})\in\mathcal{B}_{2}^{b}\left(u(\mathbf{x})\right)$, it follows from II.2 that $U(\mathbf{x}^{\prime})_{i}$ is obtained from $U(\mathbf{x})_{i}$ by two deletions. Suppose that $U(\mathbf{x})_{1}$ is $d$-regular and $\eta\left(U(\mathbf{x})_{1}\right)$ is given. According to Lemma IV.2, we can efficiently recover $U(\mathbf{x})_{1}$. Then by Lemma IV.3, the two deletions in $U(\mathbf{x})_{1}$ can be approximately located. In Lemma IV.5 below, we will show that this helps to locate errors in $u(\mathbf{x})$. To explain the idea, more notations are needed.
For $1\leq i\leq b$ and $1\leq j\leq n/b$, let $U(\mathbf{x})_{i,j}$ be the $j$-th coordinate of $U(\mathbf{x})_{i}$. By the relationship between $u(\mathbf{x})$ and $U(\mathbf{x})$, it is easy to see that $U(\mathbf{x})_{1,j}=u(\mathbf{x})_{1+(j-1)b}$. This justifies the following definition
$I_{n,b}\triangleq\left\{1+(j-1)b:1\leq j\leq n/b\right\}.$
Note that $I_{n,b}$ is a subset of $[n]$. Clearly, we have $U(\mathbf{x})_{1}=u(\mathbf{x})_{I_{n,b}}$.
Example IV.3
Let $n=10$, $b=2$ and $\mathbf{x}={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0}1{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}2}2{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}2}1{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0}1{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}1}2\in\Sigma_{3}^{10}$. Then $u(\mathbf{x})={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0}0{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}1}1{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}1}0{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0}0{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0}1$. By definition, we have $I_{n,b}=\{1,3,5,7,9\}$ and $u(\mathbf{x})_{I_{n,b}}=01100$. On the other hand, the matrix representations of $u(\mathbf{x})$ is
$U(\mathbf{x})=\begin{pmatrix}0&1&1&0&0\\ 0&1&0&0&1\end{pmatrix}$.
It is easy to see that $U(\mathbf{x})_{1}=u(\mathbf{x})_{I_{n,b}}$.
@@
Recall that $r\left(U(\mathbf{x})_{1}\right)$ is the number of runs in $U(\mathbf{x})_{1}$. For $1\leq j\leq r\left(U(\mathbf{x})_{1}\right)$, let $U(\mathbf{x})_{1,I_{j}}$ be the $j$-th run of $U(\mathbf{x})_{1}$. Furthermore, suppose $I_{j}=\left[p_{j-1}+1,p_{j}\right]$, where $p_{0}=0$, $p_{r\left(U(\mathbf{x})_{1}\right)}=n/b$ and $p_{j-1}<p_{j}$ for all $j$. By definition, $I_{j}$ is an interval in $[n/b]$. For all $j\geq 1$, we associate with $I_{j}$ an interval in $[n]$: $I_{j}^{\prime}\triangleq\left[p_{j-1}b+1,p_{j}b\right]$. In addition, define $I_{j}^{L}\triangleq\left[(p_{j-1}-1)b+2,p_{j-1}b\right]\cap[n]$. In other words, $I_{j}^{L}$ is the interval of length at most b1b-1 in $[n]$ to the left of $I_{j}^{\prime}$. Note that $I_{j}$, $I_{j}^{\prime}$ and $I_{j}^{L}$ are dependent on $\mathbf{x}$. We omit $\mathbf{x}$ for simpler notations.
Example IV.4
Let $\mathbf{x}$ be the sequence in Example IV.3. There are three runs in $U(\mathbf{x})_{1}=01100$: 0, $11$ and $00$. Recall that $p_{j}$ indicates where the $j$-th run ends. Then we have $p_{1}=1$, $p_{2}=3$, $p_{3}=5$, $I_{1}=[1,1]$, $I_{2}=[2,3]$ and $I_{3}=[4,5]$. All intervals $I_{j}^{\prime}$ and $I_{j}^{L}$ are listed in the following:
$\begin{array}[]{cc}I_{1}^{L}=\emptyset,I_{1}^{\prime}=[1,2],\\ I_{2}^{L}=[2,2],I_{2}^{\prime}=[3,6],\\ I_{3}^{L}=[6,6],I_{3}^{\prime}=[7,10].\end{array}$
We briefly explain why we define intervals $I_{j}^{\prime}$ and $I_{j}^{L}$. A $b$-burst-deletion in $u(\mathbf{x})$ induces one deletion in each row of $U(\mathbf{x})$. If it is known that the deletion in the first row $U(\mathbf{x})_{1}$ occurs in the interval $I_{j}=\left[p_{j-1}+1,p_{j}\right]$, we can only locate the deletion in the $i$-th row in the interval $\left[p_{j-1},p_{j}\right]$ when $i\geq 2$ (see II.2). Note that $U(\mathbf{x})_{1}=u(\mathbf{x})_{I_{n,b}}$. When looking at $u(\mathbf{x})$, interval $I_{j}^{L}$ (if not empty) corresponds to the last b1b-1 coordinates in the $p_{j-1}$-th column of $U(\mathbf{x})$ and interval $I_{j}^{\prime}$ corresponds to columns of $U(\mathbf{x})$ indexed by $I_{j}$. As a result, the $b$-burst-deletion in $u(\mathbf{x})$ can be located in the interval $I_{j}^{L}\cup I_{j}^{\prime}$. See Figure 1 for an illustration.
@@
In the rest of this section, let $A(\mathbf{x})=A(\mathbf{x},b)$ be the matrix representation of $\mathbf{x}$ and $A(\mathbf{x})_{1}$ be the first row of $A(\mathbf{x})$.
@@
Let $q\geq 2$, $b>1$ and $n>2b$. Suppose $\mathbf{x}\in\Sigma_{q}^{n}$ and $\mathbf{x}^{\prime}=\mathbf{x}_{[n]\setminus(D_{1}\cup D_{2})}$, where $D_{1},D_{2}\subseteq[n]$ are two disjoint intervals of length $b$. That is to say, $\mathbf{x}^{\prime}$ is obtained from $\mathbf{x}$ by a 22-$b$-burst-deletion. Suppose that $A(\mathbf{x})_{1}$ is $d$-regular, that is, $U(\mathbf{x})_{1}$ is $d$-regular. Let $\eta(\cdot)$ be the function defined in Lemma IV.2. Then given $\eta\left(U(\mathbf{x})_{1}\right)$, we can